This website uses cookies. By continuing to use this website you are agreeing to our use of cookies. 

Viewing posts from October, 2007

BADC Server Issue - Number of datasets affected
in Other
Following a technical issue with one of our servers the following datasets are not accessible at present: caesar covex dabex fluxex iasi icepic land-emiss micromix nu-wave raptor rico t-rex visurb … Read more
HadCM3-control data re-instated into archive.
in Met Office data
HadCM3-control data have been successfully restored from backup and is fully accessible once more from the BADC. Further news items regarding the server issues will be posted as progress is … Read more
BADC Server Issue - Hadcm3-control and UM data affected
in Other
Following a technical issue with one of our servers the entire HADCM3-control data and approximately 1/6th of the UM data are currently unavailable. We hope to have the HadCM3-control data … Read more
BADC FTP server Re-configuration - COMPLETED
in Other
Re-configuration of the FTP server has completed successfully. We hope that the short outages did not cause any inconvenience. Please let us know if you encounter any problems.

BADC Helpdesk …

Read more
BADC FTP server Re-configuration
in Other
Tomorrow (3rd October 2007) the BADC ftp server is being re-configured. The changes should not take long but there could be some interruption to the service throughout the day. A … Read more

Recent Posts

Archive

2023
2022
2021
2020
2019
2018
2017
2016
2015
2014
2013
2012
2011
2010
2009
2008
2007
2006
2005
2004
2003
2002
2001
2000
1996

Categories

Tags

Authors

Feeds

RSS / Atom